A given alloy contains 10% zinc and 20% copper. How many pounds of zinc and copper must be melted with 1000 pounds of the given alloy to produce another alloy analyzing 20% zinc and 24% copper? All percentages are by weight
Answer by robertb(5830)

1.00z + 0.10*1000 + 0.00c = 0.20(z + 1000 + c)
<==> z + 100 = 0.20z + 200 + 0.20c
<==> 0.80z - 0.20c = 100 <---(A)
==> 4z - c = 500 ==> c = 4z - 500
Similarly
0.00z + 0.20*1000 + 1.00c = 0.24(z + 1000 + c)
<==> 200 + c = 0.24z + 240 + 0.24c
<==> -0.24z + 0.76c = 40 <---(B)
==> -0.24z + 0.76(4z - 500) = 40
==> -0.24z + 3.04z - 380 = 40
==> 2.8z = 420
==> z = 150 lbs zinc, and
==> c = 4*150 - 500 = 100 lbs copper.